Please help, I'm sorry I'm not sure I'm in the right section. My partner is really struggling with algebra he needs to learn for an electrical installation course.
I need a clear step by step explanation on how to transpose.
I have looked as much as I can but am not sure I am looking in the right places.
Here is an example
Vp over VS = Np over Ns this transposes to Vs= Vp x Ns over Np

Could you possibly explain the rules behind transposing to me and how the equation goes from the first example to the second. I would be extremely grateful as he is pulling his hair out with it. I hope my question makes sense
Thank you

Found 2 solutions by AnlytcPhil, stanbon:
Answer by AnlytcPhil(1116)

Vp over VS = Np over Ns this 
transposes to Vs= Vp x Ns over Np

Let's do it with plain letters first.

 A     C
——— = ———
 B     D

Multiply both sides by B*D/1

B*D    A     B*D    C
——— * ——— =  ——— * ———
 1     B      1     D

Cancel the B's on the left and the D's on the right


B*D    A     B*D    C
——— * ——— =  ——— * ———
 1     B      1     D


D*A = B*C

Divide both sides by C

D*A   B*C
——— = ———
 C     C

Cancel the C's on the right:

D*A   B*C
——— = ———
 C     C

D*A   
——— = B 
 C     

or if you prefer

    D*A   
B = ———  
     C

Now put Vp where I have A, VS where I have B,
Np where I have C, and Ns where I have D.  

Edwin

Answer by stanbon(48545)

Vp over VS = Np over Ns this transposes to Vs= Vp x Ns over Np
The general rule for solving equations is "whatever you do to
one side of the equation you must do to the other side".
You have Vp/Vs = Np/Ns
You want to solve this for Vs so you need to get Vs into the numerator.
Multiply both sides by Vs to get:
Vs(Vp/Vs)=Vs(Np/Ns)
Cancel the Vs's on the left side to get:
Vp = Vs(Np/Ns)
Now multiply both sides by (Ns/Np) to get:
Vp(Ns/Np)= Vs(Np/Ns)(Ns/Np)
Cancel the Np's and the Ns's on the right side to get
Vp(Ns/Np)=Vs
This is the "transformed" equation you want.
You may think this is very cumbersome. I assure you
there are quicker ways to reach this final goal but
you wanted to see all the steps.
Cheers,
Stan H.